OpenStudy (anonymous):

f(x)=(x-2)^1/2 f'(x)=1/2(x-2)^1/2 Find equations of the tangent line(s) that pass through the point (5,2). (the point (5,2) is NOT on the graph of f.)

OpenStudy (anonymous):

Definition of the derivative was used to get f'(x). Which is the slope.

OpenStudy (ash2326):

f(x)=(x-2)^1/2 so f'(x)=1/2*(x-2)^(-1/2) we have to find equation of a tangent which passes through 5,2 let the equation be y=mx+c 2=5m+c now the tangent is touching f(x) at a,b so b=(a-2)^1/2 and the slope at this point is the slope of the tangent so slope m=1/2(a-2)^(-1/2) also (a,b) satisfy the tangent equation so b=am+c now we have 4 equations and 4 variables a, b, m and c eq.1 b=(a-2)^1/2 eq.2 b=am+c eq.3 m=1/2(a-2)^(-1/2) eq 4. 2=5m+c substitute 1 in 2, then 4 in 2 and then 3 in 2 we get \[\sqrt(a-2)=a/2*1/\sqrt(a-2)+2-5/2\sqrt(a-2)\] now put a-2=z^2 we'll get a quadratic equation in z z^2-4z+3=0 solving we get z=1,3 I proceeded with z=1 a-2=z^2 so a=3 so m=1/2*(a-2)^(-1/2) so slope =1/2 we have 2=5m+c putting value of m we get c=-1/2 so equation of tangent is y=m/2-1/2

OpenStudy (anonymous):

let m := f'(5), which is 0.5 * sqrt(3) \[y - y_0 = m(x - x_0)\]plug in 2 for y_0 and 5 for x_0 \[y = \frac{\sqrt{3}}{2}x - \frac{5\sqrt{3}}{2} + \frac{4}{2}\]

OpenStudy (ash2326):

see the graph

OpenStudy (ash2326):

you can see from graph the point where tangent touches the curve f(x) slope is 1/2

OpenStudy (anonymous):

Thank you very much! This was really helpful!

OpenStudy (ash2326):

welcome, glad to help u